OpenStudy (anonymous):

Over two summers, Ray saved $800.00 and $600.00. The polynomial 800x2 + 600x represents his savings at the beginning of the third year, where x is the growth factor. (The interest rate r is x – 1.) What is the interest rate he needs to save $1,650.00 after the third summer?

OpenStudy (anonymous):

So, Start by simplifying the original equation. 800x^2 + 600x = 1650 You get 800x^2 + 600x - 1650 = 0 16x^2 + 12x - 33 = 0 When you solve the equation, You'll get some x value, Subtract this x from 1 x-1 = your answer Let me know if you can't figure it out.

OpenStudy (anonymous):

i still don't get it

OpenStudy (anonymous):

The answer or how to solve it?

OpenStudy (anonymous):

So when you solve the polynomial (16x^2 + 12x - 33 = 0) You get x is 1.109. Now subtract 1.109 from 1 and you get----------> .11%

OpenStudy (anonymous):

Which is about the same as .1%